Please help me solve this equation:
5(x + 4)=15
The answer is supposed to be -1, but I can't get the right answer no matter what I try. This is what I have done so far:
5x + 20=15
15-20=-5
-5 X 5=-25
Your help would be greatly appreciated

Answer by fcabanski(1391)

5x + 20 = 15


5x = 15-20 = -5


Divide both sides by 5.


x = -1